Category Archives: Orbitrap status


January 23, 2015

Orbi is running Olga’s second sample, and Ninad’s two samples are next in the queue. Gaozhong’s, Ming-Yang’s, and Yang’s samples have been analyzed and the data will be processed on Monday.

I have analyzed Alex’s oligos using MALDI TOF, next will be polymers for Brooks on Monday.


January 19, 2015

Orbi was out of calibration again, so I had to recalibrate everything; it looks much better now. All gel bands have been extracted and are running now. Tomorrow, will be the negative mode day: Alex’s oligos (provided I can calibrate the Orbi).

MALDI is bored.


January 16, 2015

The instrument was not performing to the specs during the past 2 weeks, but it looks like the issue has been resolved with the help of two Thermo tech support specialists, Mike Henry and Mohit Jain. I will check the calibrations next week to make sure everything is to the specs.

Meanwhile, Orbi is analyzing Yuning’s digest. Ninad’s samples are next. I have cleaned all gel bands submitted to date, and they are currently incubating with trypsin.

MALDI-TOF is working. I ran polymer samples for Brooks. Next are oligonucleotides for Alex and milk samples for Asha.


December 22, 2014

Orbi is running second set of Scott’s samples; the first set has been analyzed and the data processed. Next in the queue are Eric’s and Josh’s samples, after which the instrument will be shut down for the holidays.

MALDI will also be shut down tomorrow to avoid potential problems (based on the last year’s holidays experience).



December 10, 2014

Orbi is running Ninad’s IPs. Eric’s samples are next. Scott’s gel lanes will be ready for proteolysis by tomorrow morning. We have heating, cooling, and electricity, all at the same time!

MALDI is experiencing severe deficit of your attention.


December 8, 2014

We are two weeks away from the winter holiday break (12/22 – 01/04), during which Orbi will be down. This morning, I received 48 samples for Orbi; so counting the 24 samples currently running, plus 6 more waiting, the Orbi queue is completely full for the 2 weeks before the break. I ask you to keep your samples safe in your refrigerators or freezers until we return in January.

MALDI currently doesn’t have a queue.


December 4, 2014

We had a power interruption today to allow electricians repair a faulty circuit breaker. All mass spectrometers were down, and now we are getting back to operating conditions. This will cause some delays in your sample analyses.

I should be able to resume running Yuanjun’s samples on Monday, after Orbi gets stable temperature so I can calibrate it.

MALDI is ready.

LCT is pumping down.

QTrap is pumping down.


December 1, 2014

Today marks my four years at Penn State.

Orbi is running John’s and Brian’s samples. Next are Ninad’s IP samples and Yuanjun’s set of 32 samples. This means Orbi will be busy until the end of this week, unless someone cuts the power again.

MALDI is bored.


November 21, 2014

We lost power to most of our equipment. OPP is here already, but we don’t know what needs to be done and whether we will be able to use our 220V power today. Orbi and MALDI-TOF are down.

Thanks to Bob Mills, we are back online. For now.

The AC has been down since this morning when we lost power, and the temperature in the lab is above 26C, so the Orbi’s LC system is down.


November 13, 2014

Orbi is running Yuanjun’s samples. The queue is approximately 40 hours including gel bands that I am currently preparing for analysis.

MALDI TOF has no queue.